OpenStudy (anonymous):

What is f(x)=x+2/6 for f(-6)? (write in simplest form) 2/3 -2/3 1/2 -1/2

OpenStudy (anonymous):

Just substitute in -6 in place of the "x" term in the function... f(x) = x + (2/6) f(-6) = -6 + (2/6) Then simplify. f(x) is like the "rule" that tells you what to do with some input, x. So f(some number) says take that number and do the rule on it. Different example: if f(x) = 2x, that means "multiply the input by 2". So f(3) would have input 3, and the output would be the input multiplied by 2... f(3) = 2(3) = 6.

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

Apparently your actual f(x) function is not x + (2/6)... none of the answers make sense with that function. It must have been f(x) = (x + 2)/6 instead. So f(6) = (-6+2)/6

OpenStudy (anonymous):

-4/6?

OpenStudy (anonymous):

Yes, good work. But you can reduce the fraction...

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

You are welcome, glad to help!